OpenStudy (anonymous):

7 - 3(x - y) = 5x + 2. What is it in standard form??

OpenStudy (ria23):

Yhu did it c: Ok, the first thing yhu want to do is distribute the -3 into the parenthesis. Do that and lemme know what yhu get

Nnesha (nnesha):

standard form mean like this Ax + By = c

OpenStudy (ria23):

Ok, so yhu equation is 7-3(x-y)=5x+2 Distributing the -3 in... Yhu get 7 - 3x + 3y = 5x + 2 correct? Now can yhu combine the like terms?

OpenStudy (anonymous):

I think I got it. 8x + 3y = 5?

OpenStudy (ria23):

c: exactly.

OpenStudy (anonymous):

Thank you(:

OpenStudy (ria23):

Wait one second...

OpenStudy (anonymous):

??

OpenStudy (ria23):

It would be 8x-3y=5

OpenStudy (ria23):

Do yhu understand why?

OpenStudy (anonymous):

Nope. I really suck at math.

OpenStudy (ria23):

Lol, yhu did really good, when yhu combine the like terms, yhu get 5+3y=8x, correct?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

Wait, how did 8x end up in the back? When they move the stuff around it confuses me.

OpenStudy (ria23):

Lol, I'll slow it down

OpenStudy (ria23):

\[7-3(x-y)=5x+2\] then yhu distribute. \[7-3x+3y=5x+2\] So yhu got it that far...

OpenStudy (ria23):

 Combine yhur like terms. Do yhu get it so far? Or did I lose yhu?